QUESTION 4 Personal Taxes Social Security Contributions Taxes on Production and Imports 20 Corporate Income Taxes Transfer Payments U.S. Exports Undistributed Corporate Profits Government Purchases Gross Private Domestic Investment 75 U.S. Imports Personal Consumption Expenditures 250 Consumption of Fixed Capital Net Foreign Factor Income Statistical Discrepancy Refer to the accompanying data (all figures in billions of dollars). Pl is $40 15 40 24 35 90 25 10 0 O A $314.00 0 B. $437.00 ° C. $346.00 $408.00
Answer #1

Ans A.) $314

PI = National income + Transfer Payments - Corporate retained earnings - corporate taxes - social security taxes.

So first we have to calculate the national income.

We calculate NI (NNP at FC) through expenditure method here.

Therefore NNPFC = personal consumption expenditure + gross domestic private investment + government purchases - consumption of fixed capital + Net exports + NFIA - Net indirect taxes (i.e taxes on production and import).

NNPFC = 250 + 75 + 90 - 25 + 2 + 10 - 20 = 382

Therefore personal income = NI + transfer payment - social security contributions - corporate income taxes - undistributed corporate profits.

so PI = 382 + 22 - 15 - 40 - 35 = $314